# LLMs in Political Science: Heralding a New Era of Visual Analysis

### v1

@@ Results
We use Gemini to annotate each image and manually evaluate Gemini annotations using a four-point Likert scale extending from 1 (poor), 2 (average), to 3 (good) and 4 (excellent) regarding their overall quality, with the understanding that a grading of 3 would indicate that Gemini annotations are potentially useful for conducting research on political images [13]. When grading differs between the authors, we assign it the lower score to be more conservative. Across the entire corpus, we observe that out of 688 images, 593 (86%) receive a rating of ‘excellent’, 49 (7%) receive a rating of ‘good’, 14 (2%) a rating of average and 32 (5%) ‘poor’. The average rating across the entire corpus is 3.8.
@@ Discussion
Image Tasks
Prompts
Gemini
Images
Face identification
List the known politicians in the image.
The image shows Nancy Pelosi, the Speaker of the United States House of Representatives.
Sentiment undertone
Use up to four words to describe the sentiment undertone of the image.
Serious, vigilant, watchful, cautious.
Image captioning
Generate a brief caption for the given image.
A group of protesters hold signs and march down a street demanding the reunification of families that were separated at the border.

### v2

@@ Results
We use Gemini to annotate each image and manually evaluate Gemini annotations using a four-point Likert scale extending from 1 (poor), 2 (average), to 3 (good) and 4 (excellent) regarding their overall quality, with the understanding that a grading of 3 would indicate that Gemini annotations are potentially useful for conducting research on political images [13]. When grading somewhat fits between two scores, we assign it the lower score to be more conservative. Across the entire corpus, we observe that out of 688 images, 593 (86%) receive a rating of ‘excellent’, 49 (7%) receive a rating of ‘good’, 14 (2%) a rating of average and 32 (5%) ‘poor’. The average rating across the entire corpus is 3.8.
@@ Discussion
Image Tasks Prompts Gemini Images
Face identification List the known politicians in the image. The image shows Nancy Pelosi, the Speaker of the United States House of Representatives.
Sentiment undertone Use up to four words to describe the sentiment undertone of the image.
Serious, vigilant, watchful, cautious. [Uncaptioned image]
Image captioning Generate a brief caption for the given image. A group of protesters hold signs and march down a street demanding the reunification of families that were separated at the border.